OpenStudy (anonymous):

I keep getting the wrong answer help :/ A barbell spins around a pivot at its center at A. The barbell consists of two small balls, each with mass m = 0.3 kg, at the ends of a massless rod of length d = 0.7 m. The barbell spins clockwise on the plane of the paper with an angular speed !0 = 20 rad/s (choice of axes: x to the right, y up, z out of the page; origin at A). Calculate the z component of the total angular momentum for this system of two particles. Answer in units of kg · m2/s

OpenStudy (anonymous):

and I seem to have pasted it on top of itself. Also sorry.

OpenStudy (anonymous):

 \[\omega = 20 rad/s=\frac{v}{r}\] ** (see the bottom for less complication) Since everything is constant, we should be able to solve at a specific time to make it easier velocity of "top" barbel at y axis is in the negative x direction \[\textbf v_{top} = -(.35m)(20rad/s) \hat{ \textbf x}=-7m/s \ \hat{ \textbf x}\] and the velocity of the "bottom" barbel at the y axis is in the positive x direction \[\textbf v_{bot} = 7m/s \ \hat{ \textbf x}\] Noting that \[\textbf r_{top} = .35m \ \hat{ \textbf y} \\ \ \\ \textbf r_{top} = -.35m \ \hat{ \textbf y}\] Then \[ \sum \textbf L = \sum_i \textbf r_i \times m_i \textbf v_i \\ \ \\ \hspace{33px}= \textbf r_{top} \times m_1 \textbf v_{top} + \textbf r_{bot} \times m_2 \textbf v_{bot} \\ \\ \hspace{33px} = m_1\big(\textbf r_{top} \times \textbf v_{top} \big) + m_2\big( \textbf r_{bot} \times \textbf v_{bot} \big) \] So you set up two cross products \[\textbf r_{top} \times \textbf v_{top} = \[\omega = 20 rad/s=\frac{v}{r}\] Since everything is constant, we should be able to solve at a specific time to make it easier velocity of "top" barbel at y axis is in the negative x direction \[\textbf v_{top} = -(.35m)(20rad/s) \hat{ \textbf x}=-7m/s \ \hat{ \textbf x}\] and the velocity of the "bottom" barbel at the y axis is in the positive x direction \[\textbf v_{bot} = 7m/s \ \hat{ \textbf x}\] Noting that \[\textbf r_{top} = .35m \ \hat{ \textbf y} \\ \ \\ \textbf r_{top} = -.35m \ \hat{ \textbf y}\] Then \[ \sum \textbf L = \sum_i \textbf r_i \times m_i \textbf v_i \\ \ \\ \hspace{33px}= \textbf r_{top} \times m_1 \textbf v_{top} + \textbf r_{bot} \times m_2 \textbf v_{bot} \\ \\ \hspace{33px} = m_1\big(\textbf r_{top} \times \textbf v_{top} \big) + m_2\big( \textbf r_{bot} \times \textbf v_{bot} \big) \] So you set up two cross products \[\textbf r_{top} \times \textbf v_{top} = \begin{vmatrix} x & y & z \\ r_x & r_y & r_z \\ v_x & v_y & v_z \end{vmatrix} = \begin{vmatrix} x & y & z \\ 0 & .35 & 0 \\ -7 & 0 & 0 \end{vmatrix}=2.45m^2/s \ \hat{ \textbf z}\] The bottom will give you the same thing, so the total angular momentum is \[ \textbf L = (m_1+m_2)2.45m^2/s \ \ \hat{ \textbf z}\] Which I totally just now realize is is exactly what you would get if you did it *not this super complicated way and just added up the two using the basic formula for moment of inertia of a point mass about an axis and angular momentum \[I = mr^2\] \[L = I \omega\] \[ \sum L = I_1\omega + I_2\omega = m_1r^2\omega + m_2r^2\omega = (m_1+m_2)(r^2\omega)\] sorry bout that.